# A model for Axial Non-Standard Interactions of neutrinos with quarks

### v1

@@ Abstract
The neutrino oscillation experiments are setting increasingly strong upper bounds on the vectorial effective Non-Standard neutrino Interactions (NSI) with matter fields. However, the bounds on the axial NSI are more relaxed, raising the hope that studying the neutral current events at an experiment such as DUNE can give a glimpse of new physics. We build a model that gives rise to axial NSI with large couplings leading to observable deviation from the standard prediction at DUNE. The model is based on a $U(1)$ gauge symmetry with a gauge boson of mass $\sim 30$ GeV which can be discovered at the high luminosity LHC. Combining the LHC and DUNE discoveries, we can unravel the axial form of interaction. The cancellation of anomalies of the gauge group suggests new heavy quarks as well as a dark matter candidate. The new quarks mixed with the first generation quarks can also be discovered at the LHC. Moreover, they provide a seesaw mechanism that explains the smallness of the $u$ and $d$ quark masses. The dark matter has an axial coupling to the quarks which makes its discovery via spin dependent direct dark matter search experiments possible.
@@ I Introduction
After establishment of the neutrino mass and mixing scheme as a solution for the solar and atmospheric neutrino anomalies, the main goal in this field has been determining the values of still unknown neutrino mass matrix parameters. In particular, to measure the Dirac CP violating phase of the PMNS neutrino mixing matrix, state-of-the-art long baseline experiments such as DUNE are under development. The precision of these experiments will enable us studying subdominant effects which might arise from new physics. Finding a footprint of new physics in the neutrino experiments is an opportunity that we do not want to miss. A wide class of new physics relevant for neutrino experiments can be described by the following effective potential
$\frac{G_{F}}{\sqrt{2}}[\bar{\nu}_{\alpha}\gamma^{\mu}(1-\gamma^{5})\nu_{\beta}][\bar{q}\gamma_{\mu}(\epsilon_{\alpha\beta}^{Vq}+\epsilon_{\alpha\beta}^{Aq}\gamma^{5})q]\ \ \ \ \ {\rm where}\ \ \ \ \ q\in\{u,d\}$ (1)
@@
This paper is organized as follows. In sect. II, with present the model and its field content and discuss the production and detection of its dark matter candidate. We also discuss bounds on the parameter space of the model. In sect. III, we discuss how various new particles introduced in this model can be discovered. In sect. IV, we first discuss generalization of the model to obtain an arbitrary lepton flavor structure for $\epsilon_{\alpha\beta}^{Au}=\epsilon_{\alpha\beta}^{Ad}$. We then suggest an alternative UV completion of the model without new colored particles. Our results are summarized in sect. V.
@@ II The model
The effective interaction of form in Eq. (1) can be obtained by integrating out a massive gauge boson, $Z^{\prime}_{\mu}$, with axial couplings to the first generation quarks as well as couplings to neutrinos. As explained before, $\epsilon_{\tau\tau}^{Aq}$ is the least experimentally constrained element. In this section, we shall introduce a model for $|\epsilon_{\tau\tau}^{Aq}|\sim 1$. Let us introduce a gauge symmetry under which all three standard quark doublets have equal charges
$Q_{1}=\left(\begin{matrix}u_{L}\cr d_{L}\end{matrix}\right)\to\exp^{i\alpha}Q_{1}\ ,\ \ \ Q_{2}=\left(\begin{matrix}c_{L}\cr s_{L}\end{matrix}\right)\to\exp^{i\alpha}Q_{2}\ ,\ \ \ {\rm and}\ \ \ \ Q_{3}=\left(\begin{matrix}t_{L}\cr b_{L}\end{matrix}\right)\to\exp^{i\alpha}Q_{3}\ .$ (2)
@@
$g_{Z}Z^{\prime}_{\mu}(\bar{u}\gamma^{\mu}\frac{1-\gamma_{5}}{2}u+\bar{d}\gamma^{\mu}\frac{1-\gamma_{5}}{2}d-\bar{u}\gamma^{\mu}\frac{1+\gamma_{5}}{2}u-\bar{d}\gamma^{\mu}\frac{1+\gamma_{5}}{2})=-g_{Z}Z^{\prime}_{\mu}(\bar{u}\gamma^{\mu}\gamma_{5}u+\bar{d}\gamma^{\mu}\gamma_{5}d)\ .$ (3)
We have therefore obtained axial form of coupling as we desired. Notice that our model respects the isospin symmetry. To cancel the $U(1)-SU(2)-SU(2)$ anomaly, we should assign $U(1)$ charges to lepton doublets, too. Since we want only $\nu_{\tau}$ to obtain effective coupling, we should assign a charge of $-9$ to $L_{3}$:
$L_{\tau}=\left(\begin{matrix}\nu_{\tau}\cr\tau_{L}\end{matrix}\right)\to\exp^{-9i\alpha}L_{\tau}\ \ \ \ {\rm under\ }U(1)\ .$ (4)
@@
$\epsilon_{\tau\tau}^{Au}=\epsilon_{\tau\tau}^{Ad}=\frac{9\sqrt{2}g_{Z}^{2}}{m_{Z^{\prime}}^{2}G_{F}}.$ (5)
@@
As mentioned above, with this $U(1)$ charge assignments to the standard left-handed fermions, the $U(1)-SU(2)-SU(2)$ anomaly will cancel out. However, we should also ensure the cancellation of the $U(1)-SU(3)-SU(3)$ anomaly. Let us denote the $U(1)$ charges of $c_{R}$, $s_{R}$, $t_{R}$ and $b_{R}$ respectively with $\alpha_{c}$, $\alpha_{s}$, $\alpha_{t}$ and $\alpha_{b}$. It is straightforward to show that in order to ensure the $U(1)-SU(3)-SU(3)$ anomaly cancellation, $\alpha_{c}+\alpha_{s}+\alpha_{t}+\alpha_{b}=8$. The measured values of the Yukawa couplings of the SM Higgs to $\tau$, $b$ and $t$ are in remarkable agreement with the SM prediction Aad et al. (2024a). Thus, it seems the mass generation mechanism for the third generation fermions to be in fact the SM Higgs mechanism. Denoting the $U(1)$ charge of the Higgs with $\alpha_{H}$, this implies $\alpha_{\tau}={-9}-\alpha_{H}$, $\alpha_{b}=1-\alpha_{H}$ and $\alpha_{t}+\alpha_{b}=2$. Throughout this paper, we shall set $\alpha_{H}=0$ so we do not need to worry about the tree level decay modes such as $H\to Z^{\prime}Z^{\prime}$. The Yukawa couplings of the second generation quarks are less constrained by observation Aad et al. (2024a) so it is imaginable that they acquire mass through coupling to another Higgs doublet with a different $U(1)$ charge. For example, if we introduce a new Higgs that couples exclusively to $s_{R}$ and $Q_{2}$, we shall have the freedom to cancel the $U(1)-SU(3)-SU(3)$ anomaly. Another way to cancel the $U(1)-SU(3)-SU(3)$ anomaly is to add new colored particles charged under the new $U(1)$. The first case without new quarks is not necessarily more economic because mass generation for the first and second generation requires four new Higgs doublet with non-trivial $U(1)$ charges. Adding new colored particles makes the model more elegant and rich from the phenomenological point of view. In the following, we shall complete our model by adding new quarks. In sect. IV, we shall discuss the case without new quarks.
@@
Let us introduce new quarks, $T_{L}$, $T_{R}$, $B_{L}$ and $B_{R}$ which are singlets of standard $SU(2)$ and are in the fundamental representation (triplet) of $SU(3)$ gauge group. The electric charges of $B$ (i.e., $B_{L}$ and $B_{R}$) and $T$ (i.e., $T_{L}$ and $T_{R}$) are respectively equal to $-1/3$ and $2/3$ as SM quarks. The field content of our model as well as their quantum numbers are shown in Table 1. The first and second generations of the leptons as well as the SM Higgs are taken to be neutral under the new $U(1)$ and are not shown in the table 1. Notice that $s_{R}$, $c_{R}$, $b_{R}$ and $t_{R}$ with charges opposite to those of $u_{R}$ and $d_{R}$ are taken to be mass eigenstates (rather than flavor eigenstates) to avoid Flavor Changing Neutral Currents (FCNC). We shall discuss more on FCNC in sect. III.
@@
Table 1: Particles and their relevant quantum number under the gauge symmetry group. Here, T, D, and S stand for Triplet, Doublet, and Singlet respectively. The first and second generation leptons are not charged under new $\rm U(1)$ gauge and are not shown here.
@@
$\rm U_{Y}(1)$ $\frac{1}{6}$ $\frac{2}{3}$ $-\frac{1}{3}$ $\frac{2}{3}$ $-\frac{1}{3}$ $-\frac{1}{2}$1 $\frac{2}{3}$ $\frac{2}{3}$ $-\frac{1}{3}$ $-\frac{1}{3}$ 00 00 00
$\rm U(1)$ $1$ $1$ $1$ $-1$ $-1$ $-9$ $-9$ $1$ $-1$ $1$ $-1$ $9$ $-2$ $-18$
With the field content shown in table 1, all gauge anomalies cancel out. The left-handed $\chi_{L}$ is added to cancel the remaining $U(1)-U(1)-U(1)$ anomaly. Being neutral, $\chi_{L}$ can be a suitable DM candidate with a dazzling prospect of direct detection as we discuss below. Notice that with this charge assignment, $u$ and $d$ quarks cannot obtain mass via Yukawa coupling to the SM Higgs. In our model, their mass comes from mixing with the fourth generation quarks. Indeed, this model also explains the hierarchy between the first and third generation quarks as a bonus in an elegant fashion. Notice that we could put $(B_{L}\ T_{L})$ as well as $(B_{R}\ T_{R})$ in the $SU(2)$ electroweak doublets but the anomaly cancellation would then require smaller $U(1)$ charge for $L_{\tau}$, leading to smaller $\epsilon_{\tau\tau}^{Au}=\epsilon_{\tau\tau}^{Ad}$.
@@
$\begin{split}&\lambda_{B}\phi\bar{B}_{L}B_{R}+\lambda_{T}\bar{T}_{L}T_{R}\phi+\frac{1}{2}\lambda_{\chi}\phi^{\prime}\chi_{L}^{T}c\chi_{L}+M_{Bd}\bar{B}_{L}d_{R}+M_{Tu}\bar{T}_{L}u_{R}+\lambda_{Bd}\bar{B}_{R}H^{\dagger}Q_{1}\\ &+\lambda_{Tu}\bar{T}_{R}H^{T}cQ_{1}+{\rm H.c.}\ \ {\rm where}\ \ c=\left(\begin{matrix}0&1\cr-1&0\end{matrix}\right).\end{split}$ (6)
We could add interactions to the second and third generation quarks as $\bar{B}_{R}H^{\dagger}Q_{2,3}$ and ${\bar{T}_{R}H^{T}c}Q_{2,3}$. For simplicity, we forbid these terms with an approximate global $U(1)$ symmetry under which only the second and third generation quarks are charged. Spontaneous symmetry breaking leads to masses for $\chi_{L}$ and the $Z^{\prime}$ boson as
$m_{\chi}=\lambda_{\chi}\langle\phi^{\prime}\rangle,\ \ {\rm and}\ \ m_{Z^{\prime}}=\sqrt{2}g_{Z^{\prime}}\left(4\langle\phi\rangle^{2}+324\langle\phi^{\prime}\rangle^{2}\right)^{1/2}$ (7)
and to the mass mixing of the first and fourth as
@@
With $M_{T}\gg M_{Tu},\lambda_{Tu}\langle H\rangle$ and $M_{B}\gg M_{Bd},\lambda_{Bd}\langle H\rangle$, there will be a seesaw mechanism that explains the smallness of the $u$ and $d$ masses relative to the electroweak scale. The mixing angles are as follows
${\rm Mixing\ between\ }u_{R}\ {\rm and}\ T_{R}=\frac{M_{Tu}^{\star}}{M_{T}}$ , ${\rm Mixing\ between\ }u_{L}\ {\rm and}\ T_{L}=\sin\alpha=\frac{\lambda_{Tu}\langle H\rangle}{M_{T}}$ (10)
@@
Through the mixing, the mass eigenstate $\tilde{T}$ and $\tilde{B}$ can decay to the first generation quarks via decay modes such as $\tilde{T}\to W+\tilde{d}$ and $\tilde{B}\to W+\tilde{u}$. Since the $U(1)$ charges of $T_{R}$ and $B_{R}$ (of $T_{L}$ and $B_{L}$) are opposite to those of $u_{R}$ and $d_{R}$ (of $Q_{1}$), in the mass basis, $Z^{\prime}$ can have flavor changing couplings to the first and fourth generation quarks with a coupling given by $2g_{Z}$ times the relevant mixing. Both decay to $W$ and to $Z^{\prime}$ are suppressed by the mixing but since we expect $g_{Z}$ to be much smaller than the $SU(2)$ coupling, the decay to the $W$ boson will dominate. In our model, since the fourth generation left-handed quarks are $SU(2)$ singlets, their electroweak couplings are different from those of the first generation. As a result, the $Z$ boson can also have flavor changing coupling between left-handed first and fourth generation quarks leading to $T\to Zu$ and $B\to Zd$. The rates of these decay modes are comparable to the charged current ones. The bound on the masses of new quarks, which mainly decay into $W$ and first generation quarks, is Aad et al. (2024b)
@@
Taking $\lambda_{T},\lambda_{B}$ in the perturbative range $(\lambda_{T},\lambda_{B}<4)$, Eqs. (5,7) yield
@@
As a result, our model can lead to $\epsilon_{\tau\tau}^{Au}=\epsilon_{\tau\tau}^{Ad}>0.1$ which can be probed by DUNE Abbaslu et al. (2024). Fig. 1 shows $\epsilon_{\tau\tau}^{Au}=\epsilon_{\tau\tau}^{Ad}$ versus $m_{\chi}$ for different values of Yukawa coupling fixing $M_{T}$ and $M_{B}$ to 1550 GeV. For $m_{\chi}<10$ GeV, $\langle\phi\rangle\ll M_{T}/\lambda_{T},M_{B}/\lambda_{B}$ so the NSI coupling becomes independent of $m_{\chi}$. Notice that saturating the bound in Eq. (15) requires $9\langle\phi^{\prime}\rangle\ll\langle\phi\rangle$ (and therefore $m_{\chi}\ll M_{T},M_{B}$). As we shall see below, this is the range that we are interested in. Taking $M_{Z^{\prime}}>30$ GeV, we can safely use the four fermion effective interaction for neutrino scattering in experiments such as DUNE. Such large values of $m_{Z^{\prime}}$ requires $g_{Z}>0.02.$ On the other hand, to avoid the present bound Sirunyan et al. (2019), $g_{Z}$ should be smaller than 0.2.

### v2

@@ Abstract
The neutrino oscillation experiments are setting increasingly strong upper bounds on the vector Non-Standard neutrino Interactions (NSI) with matter fields. However, the bounds on the axial NSI are more relaxed, raising the hope that studying the neutral current events at an experiment such as DUNE can give a glimpse on new physics. We build a model that gives rise to axial NSI with large couplings leading to observable deviation from the standard prediction at DUNE. The model is based on a $U(1)$ gauge symmetry with a gauge boson of mass $\sim 30$ GeV which can be discovered at the high luminosity LHC. Combining the LHC and DUNE discoveries, we can unravel the axial form of interaction. The cancellation of anomalies of the gauge group suggests new heavy quarks as well as a dark matter candidate. The new quarks mixed with the first generation quarks can also be discovered at the LHC. Moreover, they provide a seesaw mechanism that explains the smallness of the $u$ and $d$ quark masses. The dark matter has an axial coupling to the quarks which makes its discovery via spin dependent direct dark matter search experiments possible.
@@ I Introduction
After the establishment of the neutrino mass and mixing scheme as a solution for the solar and atmospheric neutrino anomalies, the main goal in this field has been determining the values of still unknown neutrino mass matrix parameters. In particular, to measure the Dirac CP violating phase of the PMNS neutrino mixing matrix, state-of-the-art long baseline experiments such as DUNE are under development. The precision of these experiments will enable us to study subdominant effects that might arise from new physics. Finding a footprint of new physics in the neutrino experiments is an opportunity that we do not want to miss. A wide class of new physics relevant for neutrino experiments can be described by the following effective potential
$\frac{G_{F}}{\sqrt{2}}[\bar{\nu}_{\alpha}\gamma^{\mu}(1-\gamma^{5})\nu_{\beta}][\bar{q}\gamma_{\mu}(\epsilon_{\alpha\beta}^{Vq}+\epsilon_{\alpha\beta}^{Aq}\gamma^{5})q]\ \ \ \ \ {\rm where}\ \ \ \ \ q\in\{u,d\}\ .$ (1)
@@
$0.19<\epsilon_{ee}^{Au}-\epsilon_{ee}^{Ad}<0.13\ \ {\rm and}\ \ -0.13<\epsilon_{e\tau}^{Au}-\epsilon_{e\tau}^{Ad}<0.1\ .$
Moreover, CHARM implies
$1.7<\epsilon_{ee}^{Au}<0.7,\ -0.8<\epsilon_{ee}^{Ad}<0.9\ {\rm and}\ |\epsilon_{e\tau}^{Aq}|<1\ .$
The strongest bounds on the $\mu\mu$ and $\mu\tau$ components come from the NuTeV neutrino scattering experiment Zeller et al. (2002):
$|\epsilon^{Aq}_{\mu\mu}|<0.01\ \ {\rm and}\ \ |\epsilon^{Aq}_{\mu\tau}|<0.1\ .$
The $e\mu$ components are severely constrained by the bound on the lepton flavor violating process $\mu^{-}{\rm Ti}\to e^{-}{\rm Ti}$ Davidson et al. (2003):
$|\epsilon_{e\mu}^{Aq}|<7.7\times 10^{-4}\ .$
The SNO experiment Coloma et al. (2023) finds two solutions for the difference of $\epsilon_{\tau\tau}^{Ad}$ and $\epsilon_{\tau\tau}^{Au}$ as
$2.1<\epsilon_{\tau\tau}^{Au}-\epsilon_{\tau\tau}^{Ad}<-1.8\ \ {\rm or}\ \ -0.2<\epsilon_{\tau\tau}^{Au}-\epsilon_{\tau\tau}^{Ad}<0.15\ .$
Despite these bounds, the $\tau\tau$ component of the isospin singlet combination, $\epsilon^{Au}+\epsilon^{Ad}$, is still practically unconstrained in the present literature, allowing values even larger than 1 Coloma et al. (2023). [^1] Studying the neutral current at MINOS and MINOS+ can also constrain $\epsilon_{\tau\tau}^{Au}$ and $\epsilon_{\tau\tau}^{Ad}$. Recently, we have set a bound $0.35<\epsilon_{\tau\tau}^{Au}=\epsilon_{\tau\tau}^{Ad}<0.3$ Abbaslu and Farzan (2025) using the MINOS and MINOS+ data. Moreover in Abbaslu et al. (2025), we discuss the possibility of constraining $\epsilon_{\tau\tau}^{Aq}$ by the KamLAND data. Ref. Abbaslu et al. (2024) shows that studying the neutral current events at the near and far detectors of DUNE can significantly improve the information on the axial NSI. In particular, $\epsilon^{Au}_{\tau\tau}$ and $\epsilon^{Ad}_{\tau\tau}$ can be probed down to $O(0.1)$ Abbaslu et al. (2024).
@@
This paper is organized as follows. In sect. II, we present the model and its field content and discuss the production and detection of its dark matter candidate. We also discuss the bounds on the parameter space of the model. In sect. III, we discuss how various new particles introduced in this model can be discovered. In sect. IV, we first discuss a generalization of the model to obtain an arbitrary lepton flavor structure for $\epsilon_{\alpha\beta}^{Au}=\epsilon_{\alpha\beta}^{Ad}$. We then suggest an alternative UV completion of the model without new colored particles. Our results are summarized in sect. V.
@@ II The model
The effective interaction of the form in Eq. (1) can be obtained by integrating out a massive gauge boson, $Z^{\prime}_{\mu}$, with axial couplings to the first generation quarks as well as with couplings to neutrinos. As explained before, $\epsilon_{\tau\tau}^{Aq}$ is the least experimentally constrained element. In this section, we shall introduce a model for $|\epsilon_{\tau\tau}^{Aq}|\sim 1$. Let us introduce a gauge symmetry under which all three standard quark doublets have equal charges
$Q_{1}=\left(\begin{matrix}u_{L}\cr d_{L}\end{matrix}\right)\to e^{i\alpha}Q_{1}\ ,\ \ \ Q_{2}=\left(\begin{matrix}c_{L}\cr s_{L}\end{matrix}\right)\to e^{i\alpha}Q_{2}\ ,\ \ \ {\rm and}\ \ \ \ Q_{3}=\left(\begin{matrix}t_{L}\cr b_{L}\end{matrix}\right)\to e^{i\alpha}Q_{3}\ .$ (2)
@@
$\begin{split}&g_{Z^{\prime}}Z^{\prime}_{\mu}\left(\bar{u}\gamma^{\mu}\frac{1-\gamma_{5}}{2}u+\bar{d}\gamma^{\mu}\frac{1-\gamma_{5}}{2}d-\bar{u}\gamma^{\mu}\frac{1+\gamma_{5}}{2}u-\bar{d}\gamma^{\mu}\frac{1+\gamma_{5}}{2}d\right)\\ &=-g_{Z^{\prime}}Z^{\prime}_{\mu}\left(\bar{u}\gamma^{\mu}\gamma_{5}u+\bar{d}\gamma^{\mu}\gamma_{5}d\right)\ .\end{split}$ (3)
We have therefore obtained the axial form of coupling as we desired. Notice that our model respects the isospin symmetry. To cancel the U(1)SU(2)SU(2)U(1)-SU(2)-SU(2) anomaly, we should assign $U(1)$ charges to lepton doublets, too. Since we want only $\nu_{\tau}$ to obtain effective coupling, we should assign a charge of 9-9 to $L_{3}$:
$L_{\tau}=\left(\begin{matrix}\nu_{\tau}\cr\tau_{L}\end{matrix}\right)\to e^{-9i\alpha}L_{\tau}\ \ \ \ {\rm under\ }U(1)\ .$ (4)
@@
$\epsilon_{\tau\tau}^{Au}=\epsilon_{\tau\tau}^{Ad}=\frac{9\sqrt{2}g_{Z^{\prime}}^{2}}{m_{Z^{\prime}}^{2}G_{F}}.$ (5)
As mentioned above, with this $U(1)$ charge assignments to the standard left-handed fermions, the U(1)SU(2)SU(2)U(1)-SU(2)-SU(2) anomaly will cancel out. However, we should also ensure the cancellation of the U(1)SU(3)SU(3)U(1)-SU(3)-SU(3) anomaly.
@@
Before introducing our mechanism for the cancellation of the U(1)SU(3)SU(3)U(1)-SU(3)-SU(3) anomaly, let us discuss the $U(1)$ charges of the right-handed fermions as well as that of the Higgs. We denote the $U(1)$ charges of $\tau_{R}$ $c_{R}$, $s_{R}$, $t_{R}$ and $b_{R}$ respectively by $\alpha_{\tau}$, $\alpha_{c}$, $\alpha_{s}$, $\alpha_{t}$ and $\alpha_{b}$. The measured values of the Yukawa couplings of the SM Higgs to $\tau$, $b$ and $t$ are in remarkable agreement with the SM prediction Aad et al. (2024a). Thus, it is established that the mass generation mechanism for the third generation fermions is indeed the SM Higgs mechanism. Denoting the $U(1)$ charge of the Higgs by $\alpha_{H}$, this implies $\alpha_{\tau}={-9}-\alpha_{H}$, $\alpha_{b}=1-\alpha_{H}$ and $\alpha_{t}+\alpha_{b}=2$. Throughout this paper, we shall set $\alpha_{H}=0$ so we do not need to worry about the tree level decay modes such as $H\to Z^{\prime}Z^{\prime}$. The Yukawa couplings of the second generation quarks are less constrained by observation Aad et al. (2024a) so it is imaginable that they acquire mass through couplings to another Higgs doublet with a $U(1)$ charge of $\alpha_{H^{\prime}}$. For example, if we introduce a new Higgs that couples exclusively to $s_{R}$ and $Q_{2}$, the $U(1)$ charge cancelation implies $\alpha_{s}+\alpha_{H^{\prime}}=\alpha$ so, with an arbitrary value of $\alpha_{H^{\prime}}$, the mass generation for the $s$ quark does not dictate the equality of $\alpha_{s}$ and $\alpha$. Taking $\alpha_{s}=8-\alpha_{c}-\alpha_{t}-\alpha_{b}$, the U(1)SU(3)SU(3)U(1)-SU(3)-SU(3) anomaly cancels out without needing new quarks. However, the mass generation for the first and second generations requires four new Higgs doublets with non-trivial $U(1)$ charges. In sect. IV, we shall discuss the variant of our model without new quarks.
@@
Another way to cancel the U(1)SU(3)SU(3)U(1)-SU(3)-SU(3) anomaly is to add new colored particles charged under the new $U(1)$. Adding new colored particles makes the model more elegant and rich from the phenomenological point of view. In the following, we shall complete our model by adding new quarks. Let us introduce new quarks, $T_{L}$, $T_{R}$, $B_{L}$ and $B_{R}$ which are singlets of standard $SU(2)$ and are in the fundamental representation (triplet) of the $SU(3)$ gauge group. The electric charges of $B$ (i.e., $B_{L}$ and $B_{R}$) and $T$ (i.e., $T_{L}$ and $T_{R}$) are respectively equal to 1/3-1/3 and $2/3$ as SM quarks. The field content of our model as well as their quantum numbers are shown in Table 1. The first and second generations of the leptons as well as the SM Higgs are taken to be neutral under the new $U(1)$ and are not shown in the table 1. Notice that $s_{R}$, $c_{R}$, $b_{R}$, and $t_{R}$ with charges opposite to those of $u_{R}$ and $d_{R}$ are taken to be mass eigenstates (rather than flavor eigenstates) to avoid Flavor Changing Neutral Currents (FCNC). We shall discuss more on FCNC in sect. III.
@@
Table 1: Particles and their relevant quantum number under the gauge symmetry group. Here, T, D, and S stand for Triplet, Doublet, and Singlet respectively. The first and second generation leptons are not charged under the new $\rm U(1)$ gauge and are not shown here.
@@
$\rm U_{Y}(1)$ $\frac{1}{6}$ $\frac{2}{3}$ $\frac{1}{3}$ $\frac{2}{3}$ $\frac{1}{3}$ $\frac{1}{2} -$1 $\frac{2}{3}$ $\frac{2}{3}$ $\frac{1}{3}$ $\frac{1}{3}$ 0 0 0
$\rm U(1)$ 11 11 11 1-1 1-1 9-9 9-9 11 1-1 11 1-1 99 2-2 18-18
With the field content shown in table 1, all gauge anomalies cancel out. The left-handed $\chi_{L}$ is added to cancel the remaining $U(1)U(1)$-U(1)-U(1) anomaly. Being neutral, $\chi_{L}$ can be a suitable DM candidate with a dazzling prospect of direct detection as we discuss below. Notice that with this charge assignment, the $u$ and $d$ quarks cannot obtain mass via Yukawa coupling to the SM Higgs. In our model, their masses come from mixing with the fourth generation quarks. Indeed, this model also explains the hierarchy between the first and third generation quarks as a bonus in an elegant fashion. Notice that we could put $(B_{L}\ T_{L})$ as well as $(B_{R}\ T_{R})$ in the $SU(2)$ electroweak doublets but the anomaly cancellation would then require smaller $U(1)$ charge for $L_{\tau}$, leading to smaller $\epsilon_{\tau\tau}^{Au}=\epsilon_{\tau\tau}^{Ad}$.
@@
$\begin{split}&\lambda_{B}\phi\bar{B}_{L}B_{R}+\lambda_{T}\bar{T}_{L}T_{R}\phi+\frac{1}{2}\lambda_{\chi}\phi^{\prime}\chi_{L}^{T}c\chi_{L}+M_{Bd}\bar{B}_{L}d_{R}+M_{Tu}\bar{T}_{L}u_{R}+\lambda_{Bd}\bar{B}_{R}H^{\dagger}Q_{1}\\ &+\lambda_{Tu}\epsilon_{ij}H_{i}\bar{T}_{R}(Q_{1})_{j}+{\rm H.c.}\ \ {\rm where}\ \ c=\left(\begin{matrix}0&1\cr-1&0\end{matrix}\right).\end{split}$ (6)
Like the matrix $c$, $\epsilon_{ij}$ is also an anti-symmetric matrix with off-diagonal elements equal to $\pm 1$. While the $\epsilon$ matrix acts on the $SU(2)$ indices, the $c$ matrix acts on the spinorial indices of the Weyl fermions. We could add interactions to the second and third generation quarks as $\bar{B}_{R}H^{\dagger}Q_{2,3}$ and $\epsilon_{ij}H_{i}\bar{T}_{R}(Q_{2,3})_{j}$. For simplicity, we forbid these terms with an approximate global $U(1)$ symmetry under which only the second and third generation quarks are charged. Spontaneous symmetry breaking leads to a Majorana mass for $\chi_{L}$
$\frac{1}{2}m_{\chi}\chi_{L}^{T}c\chi_{L}\ \ \ {\rm where}\ \ \ m_{\chi}=\lambda_{\chi}\langle\phi^{\prime}\rangle\ .\ \ \$(7)
Notice that $\chi_{L}^{T}c=\bar{\chi}_{L}^{c}$. For a detailed discussion of the CP-conjugation of the Majorana fields, the readers may consult Ref. Fujikawa and Tureanu (2024). Moreover, vacuum expectation values of $\phi$ and $\phi^{\prime}$ lead to a mass for the gauge boson
$m_{Z^{\prime}}=\sqrt{2}g_{Z^{\prime}}\left(4\langle\phi\rangle^{2}+324\langle\phi^{\prime}\rangle^{2}\right)^{1/2}\ .$ (8)
After the $U(1)$ and $SU(2)\times U_{Y}(1)$ spontaneous symmetry breaking, the first and fourth generation quarks obtain mass and mix as
@@
Relations in Eq. (10) are written at the energy scale around $M_{B}$ and $M_{T}$. To compare with the so-called “current quark masses” reported in PDG Navas et al. (2024) at the energy scale $\sim 2$ GeV, we should take into account the running of the quark masses with energy which results in an enhancement of order of one Chetyrkin and Retey (2000). Notice that the running of quark masses from the mass scale of $T$ and $B$ down to 2 GeV is similar to SM because the new quarks decouple below their mass scale. The coupling of the new gauge interaction is much smaller than the QCD coupling and its contributions to the running of the masses are therefore negligible. In our model, there is no upper bound on the mass parameters $M_{Bd}$ and $M_{Tu}$. However, their smallness relative to other mass scales of the model is “technically natural.” That is at the limit that $M_{Bd},M_{Tu}\to 0$, a global symmetry under which $\phi\to e^{i\alpha}\phi$, $T_{L}\to e^{i\alpha}T_{L}$ and $B_{L}\to e^{i\alpha}B_{L}$ is restored, satisfying the famous ’t Hooft criterion for naturalness Dine (2015). Notice that $M_{Bd}$ and $M_{Tu}$ are dimensionful parameters so the breaking of the global symmetry is soft which implies that the radiative corrections to $M_{Bd}$ and $M_{Tu}$ are proportional to themselves. Below the mass scale of $T$ and $B$, these new quarks decouple so the running of $M_{Bd}$ and $M_{Tu}$ below this scale is of no relevance. With $M_{T}\gg M_{Tu},\lambda_{Tu}\langle H\rangle$ and $M_{B}\gg M_{Bd},\lambda_{Bd}\langle H\rangle$, there will be a seesaw mechanism that explains the smallness of the $u$ and $d$ masses relative to the electroweak scale. The mixing angles are as follows
${\rm Mixing\ between\ }u_{R}\ {\rm and}\ T_{R}=\frac{M_{Tu}^{\star}}{M_{T}}$ , ${\rm Mixing\ between\ }u_{L}\ {\rm and}\ T_{L}=\sin\alpha=\frac{\lambda_{Tu}\langle H\rangle}{M_{T}},$ (11)
@@
Through the mixing, the mass eigenstate $\tilde{T}$ and $\tilde{B}$ can decay to the first generation quarks via decay modes such as $\tilde{T}\to W+\tilde{d}$ and $\tilde{B}\to W+\tilde{u}$. Since the $U(1)$ charges of $T_{R}$ and $B_{R}$ (of $T_{L}$ and $B_{L}$) are opposite to those of $u_{R}$ and $d_{R}$ (of $Q_{1}$), in the mass basis, $Z^{\prime}$ can have flavor changing couplings to the first and fourth generation quarks with a coupling given by $2g_{Z^{\prime}}$ times the relevant mixing. Both decays to $W$ and to $Z^{\prime}$ are suppressed by the mixing but since we expect $g_{Z^{\prime}}$ to be much smaller than the $SU(2)$ coupling, the decay to the $W$ boson will dominate. In our model, since the fourth generation left-handed quarks are $SU(2)$ singlets, their electroweak couplings are different from those of the first generation. As a result, the $Z$ boson can also have flavor changing coupling between left-handed first and fourth generation quarks leading to $T\to Zu$ and $B\to Zd$. The rates of these decay modes are comparable to the charged current ones. The bound on the masses of new quarks, which mainly decay into $W$ and first generation quarks, is Aad et al. (2024b)
@@
Taking $\lambda_{T}$ and $\lambda_{B}$ in the perturbative range $(\lambda_{T},\lambda_{B}<4)$, Eqs. (5,8) yield
@@
As a result, our model can lead to $\epsilon_{\tau\tau}^{Au}=\epsilon_{\tau\tau}^{Ad}>0.1$ which can be probed by DUNE Abbaslu et al. (2024). Notice that with large $\lambda_{T}$ and $\lambda_{B}$, the radiative corrections above the $T$ and $B$ masses can be significant; however, as long as the loop suppression factor, $N_{c}\lambda_{T}^{2}/(16\pi^{2})$ or $N_{c}\lambda_{B}^{2}/(16\pi^{2})$ (with $N_{c}=$number of colors=3) is smaller than 1, the model stays in the perturbative regime Ishiwata and Wise (2011); Allwicher et al. (2021). Figure 1 shows $\epsilon_{\tau\tau}^{Au}=\epsilon_{\tau\tau}^{Ad}$ versus $m_{\chi}$ for different values of the Yukawa couplings, fixing $M_{T}$ and $M_{B}$ to 1550 GeV. For $m_{\chi}<10$ GeV, $\langle\phi\rangle\ll M_{T}/\lambda_{T},M_{B}/\lambda_{B}$ so the NSI coupling becomes independent of $m_{\chi}$. Notice that saturating the bound in Eq. (16) requires $9\langle\phi^{\prime}\rangle\ll\langle\phi\rangle$ (and therefore $m_{\chi}\ll M_{T},M_{B}$). As we shall see below, this is the range that we are interested in. Taking $M_{Z^{\prime}}>30$ GeV, we can safely use the four fermion effective interaction for neutrino scattering in experiments such as DUNE. Such large values of $m_{Z^{\prime}}$ requires $g_{Z^{\prime}}>0.02.$ On the other hand, to avoid the present bound Sirunyan et al. (2019), $g_{Z^{\prime}}$ should be smaller than 0.2.